OpenStudy (anonymous):

WILL FAN AND MEDAL Jerome has 1/4 of groups video games at his house. Mario has 2/5 of the groups video games at his house. What fraction of the groups video games is either at Jeromes house or Marios house? also plz explain when u find the answer

OpenStudy (anonymous):

HELP PLZ

OpenStudy (anonymous):

1 4. 1 ---- x ---- =. ------ 4. 5 5 You have to multiply in order to know how many video games was in their houses. If you subtract it then you would have got the wiring answer.

OpenStudy (anonymous):

I meant: 1. 4. 1 ----- x. ----- =. ------- 4. 5. 5

OpenStudy (anonymous):

THANK YOU SO MUCH

OpenStudy (anonymous):

\[1-\frac{1}{4}-\frac{2}{5}=\frac{7}{20} \]

OpenStudy (anonymous):

I have another question Rebecca had 9/10 of a yard of ribbon. Then she gave 7/20 of yard to a friend. How much does she have left explain